# Generalization Error of the Tilted Empirical Risk

### v1

@@ 4.1 Uniform Bounds
Under the same assumptions in Theorem 4.6 and a finite hypothesis space, then for $n\geq\frac{(4\gamma^{2}M_{2}^{u}+8/3\zeta)\log(2/\delta)}{\zeta^{2}\exp(2\gamma R_{T}^{u})}$ with probability at least $(1-\delta)$ and $\gamma<0$, the excess risk of tilted empirical risk satisfies
$\begin{split}\mathfrak{E}_{\gamma}(\mu)&\leq\frac{4\exp(-\gamma R_{T}^{u})}{(1-\zeta)}\sqrt{\frac{M_{2}^{u}(\log(\mathrm{card}(\mathcal{H}))+\log(2/\delta))}{n}}\\ &\qquad-\frac{8\exp(-\gamma R_{T}^{u})(\log(\mathrm{card}(\mathcal{H}))+\log(2/\delta))}{3n\gamma(1-\zeta)}-\gamma M_{2}^{u}.\end{split}$ (23)
4.2 Information-theoretic Bounds
Proposition 4.10.
Given Assumption 4.3 and Assumption 4.4, the following inequality holds for $\gamma<0$,
$\begin{split}\mathrm{R}_{\gamma}(H,P_{H}\otimes\mu)-\mathrm{R}_{\gamma}(H,P_{H,S})\leq\begin{cases}\exp(-\gamma R_{T})\sqrt{\frac{M_{2}I(H;S)}{n}},\qquad&\mbox{ if }\frac{I(H;S)}{n}\leq\frac{\gamma^{2}M_{2}}{2}\\ \frac{-\exp(-\gamma R_{T})}{\gamma}\Big{(}\frac{I(H;S)}{n}+\frac{\gamma^{2}M_{2}}{2}\Big{)},\qquad&\mbox{ if }\frac{I(H;S)}{n}>\frac{\gamma^{2}M_{2}}{2}\end{cases}\end{split}$
Corollary 4.11.
Given Assumption 4.4 and Assumption 4.3, the following inequality holds for $\gamma<0$,
$\begin{split}\mathrm{R}_{\gamma}(H,P_{H}\otimes\mu)-\mathrm{R}_{\gamma}(H,P_{H,S})\geq\begin{cases}-\exp(-\gamma R_{T})\sqrt{\frac{M_{2}I(H;S)}{n}},\quad&\mbox{ if }\max\big{(}\frac{2I(H;S)}{\gamma^{2}M_{2}},\frac{\gamma^{2}M_{2}I(H;S)}{\exp(2\gamma R_{T})}\big{)}\leq n\\ \frac{\exp(-\gamma R_{T})}{\gamma}\Big{(}\frac{I(H;S)}{n}+\frac{\gamma^{2}M_{2}}{2}\Big{)},\quad&\mbox{ if }\min(\frac{\exp(\gamma R_{T})-\frac{\gamma^{2}M_{2}}{2}}{I(H;S)},\frac{2I(H;S)}{\gamma^{2}M_{2}})>n.\end{cases}\end{split}$
Using Proposition 4.10, we can obtain the following upper bound on the expected generalization error.
Theorem 4.12.
Given Assumption 4.3 and Assumption 4.4, the following upper bound holds on the expected tilted generalization error for $\gamma<0$,
$\begin{split}\overline{\mathrm{gen}}_{\gamma}(H,S)&\leq\begin{cases}\exp(-\gamma R_{T})\sqrt{\frac{M_{2}I(H;S)}{n}}-\frac{\gamma}{2}M_{2},\qquad&\mbox{ if }\frac{I(H;S)}{n}\leq\frac{\gamma^{2}M_{2}}{2}\\ \frac{-\exp(-\gamma R_{T})}{\gamma}\Big{(}\frac{I(H;S)}{n}+\frac{\gamma^{2}M_{2}}{2}\Big{)}-\frac{\gamma}{2}M_{2},\qquad&\mbox{ if }\frac{I(H;S)}{n}>\frac{\gamma^{2}M_{2}}{2}\end{cases}\end{split}$
Remark 4.13.
Assuming $\gamma=O(n^{-1/2})$, the upper bound in Theorem 4.12 has the convergence rate $O(n^{-1/2})$. Note that the result in Theorem 4.12 holds for unbounded loss functions, provided that the second moment of the loss function exists.
A similar approach to Theorem 4.12 can be applied to derive an information-theoretical lower bound on the expected tilted generalization error for $\gamma<0$ by applying Corollary 4.11.
Theorem 4.14.
Given Assumption 4.3 and Assumption 4.4, the following lower bound holds on the expected tilted generalization error for $\gamma<0$,
$\begin{split}\overline{\mathrm{gen}}_{\gamma}(H,S)&\geq\begin{cases}-\exp(-\gamma R_{T})\sqrt{\frac{M_{2}I(H;S)}{n}}+\frac{\gamma}{2}M_{2},\qquad&\mbox{ if }\max\big{(}\frac{2I(H;S)}{\gamma^{2}M_{2}},\frac{\gamma^{2}M_{2}I(H;S)}{\exp(2\gamma R_{T})}\big{)}\leq n\\ \frac{\exp(-\gamma R_{T})}{\gamma}\Big{(}\frac{I(H;S)}{n}+\frac{\gamma^{2}M_{2}}{2}\Big{)}+\frac{\gamma}{2}M_{2},\qquad&\mbox{ if }\min(\frac{\exp(\gamma R_{T})-\frac{\gamma^{2}M_{2}}{2}}{I(H;S)},\frac{2I(H;S)}{\gamma^{2}M_{2}})>n.\end{cases}\end{split}$
@@ 5 Robustness of TERM
As shown in experiments by Li et al. (2021), the tilted empirical risk is robust to noise or outliers samples during training using negative tilt ($\gamma<0$). In this section, we study the robustness of the TER under distributional shift, $\tilde{\mu}$ and the following assumptions are made.
Assumption 5.1 (Uniform bounded true risk under $\tilde{\mu}$).
There is a constant $R_{T}^{s}$ such that the loss function $\ell$ satisfies $\mathbb{E}_{\tilde{\mu}}[\ell(h,\tilde{Z})]\leq R_{T}^{s}<\infty$ uniformly for all $h\in\mathcal{H}$.
@@
There is a constant $M_{2}^{s}$ such that the loss function $(H,Z)\mapsto\ell(H,Z)$ satisfies $\mathbb{E}_{\tilde{\mu}}[\ell^{2}(h,Z)]\leq M_{2}^{s}$ uniformly for all $h\in\mathcal{H}$.
@@
Given Assumption 4.1, then the difference of tilted population risk under, (9), between $\mu$ and $\tilde{\mu}$ is bounded as follows,
$\begin{split}&\frac{1}{\gamma}\log(\mathbb{E}_{\tilde{Z}\sim\mu}[\exp(\gamma\ell(h,\tilde{Z}))])-\frac{1}{\gamma}\log(\mathbb{E}_{\tilde{Z}\sim\tilde{\mu}}[\exp(\gamma\ell(h,\tilde{Z}))])\leq\frac{\mathbb{TV}(\mu,\tilde{\mu})}{|\gamma|\exp(\gamma R_{T}^{u})}.\end{split}$ (24)
Note that, for positive $\gamma$, the result in Proposition 5.3 does not hold and can be unbounded. Using Proposition 5.3, we can provide an upper bound on the tilted generalization error under distributional shift.
@@
Given Assumptions 4.1, 4.2, 5.1 and 5.2, for any fixed $h\in\mathcal{H}$ and with probability least $(1-\delta)$ for $\gamma<0$, then the following upper bound holds on the tilted generalization error
$\begin{split}\mathrm{gen}_{\gamma}(h,\hat{S})&\leq 2\exp(-\gamma R_{T}^{s})\sqrt{\frac{M_{2}^{s}(\log(2/\delta))}{n}}\\ &\qquad-\frac{4\exp(-\gamma R_{T}^{s})(\log(2/\delta))}{3n\gamma}-\frac{\gamma}{2}M_{2}^{u}-\frac{\exp(-\gamma R_{T}^{u})\mathbb{TV}(\mu,\tilde{\mu})}{\gamma},\end{split}$
where $\hat{S}$ is the training dataset under the distributional shift.
It is noteworthy that the upper bound in Theorem 5.4 can be infinite for $\gamma\rightarrow-\infty$ and $\gamma=0$. Consequently, there must exist a $\gamma\in(-\infty,0)$ that minimizes this upper bound. To illustrate this point, consider the case where $n\rightarrow\infty$; here, the first and second terms in the upper bound would vanish. Thus, we are led to the following minimization problem:
$\gamma^{\star}:=\mathop{\mathrm{arg},\min}_{\gamma\in(-\infty,0)}\left[-\frac{\gamma}{2}M_{2}^{u}-\frac{\exp(-\gamma R_{T}^{u})\mathbb{TV}(\mu,\tilde{\mu})}{\gamma}\right],$ (25)
Similar results to Theorem 5.4, can be derived via an information-theoretic approach.
Robustness vs Generalization: The term $\frac{\mathbb{TV}(\mu,\tilde{\mu})}{|\gamma|\exp(\gamma R_{T}^{u})}$ represents the distributional shift cost (or robustness) associated with the TER. This cost can be reduced by increasing $|\gamma|$. However, increasing $|\gamma|$ also amplifies other terms in the upper bound of the tilted generalization error. Therefore, there is a trade-off between robustness and generalization, particularly for $\gamma<0$ in the TER. Interestingly, Li et al. (2021) also observed this trade-off for negative tilt.
@@ 6 The KL-Regularized TERM Problem
where $I(H;S)\leq\mathrm{KL}(P_{H|S}\|\pi_{H}|P_{S})$ and $\pi_{H}$ is a prior distribution over hypothesis space $\mathcal{H}$. All proof details are deferred to Appendix F.
@@
The solution to the expected TERM regularized via KL divergence, (27), is the tilted Gibbs Posterior (algorithm),
@@
is the solution to the KL-regularized ERM minimization problem, where $\tilde{F}_{\alpha}(S)$ is the normalization factor. Therefore, the tilted Gibbs posterior is different from the Gibbs posterior, (29). It can be shown that for $\gamma\rightarrow 0$, the tilted Gibbs posterior converges to the Gibbs posterior. Therefore, it is interesting to study the expected generalization error of the tilted Gibbs posterior. For this purpose, we give an exact characterization of the difference between the expected TER under the joint and the product of marginal distributions of $H$ and $S$. More results regarding the Gibbs posterior are provided in Appendix F.
@@
$\overline{\mathrm{gen}}_{\gamma}(H,S)\leq\frac{\alpha(\exp(|\gamma|M)-1)^{2}}{2\gamma^{2}n}-\frac{\gamma\exp(-\gamma M)}{2}\big{(}1-\frac{1}{n}\big{)}\mathbb{E}_{P_{H}}\big{[}\mathrm{Var}_{\tilde{Z}\sim\mu}(\ell(H,\tilde{Z}))\big{]}.$ (31)
@@
$|\overline{\mathrm{gen}}_{\gamma}(H,S)|\leq\frac{\alpha(\exp(|\gamma|M)-1)^{2}}{2\gamma^{2}n}+\frac{|\gamma|M^{2}\exp(|\gamma|M)}{8}\Big{(}1-\frac{1}{n}\Big{)}.$ (32)
@@ 7 Related Works
Generalization Error Analysis: Different approaches have been applied to study the generalization error of general learning problems under empirical risk minimization, including VC dimension-based, Rademacher complexity, PAC-Bayesian, stability and information-theoretic bounds. In this section, we discuss the related works about Uniform and information-theoretic bounds. More related works for generalization error analysis are discussed in Appendix A.
@@
The aforementioned approaches are applied to study the generalization error in the linear empirical risk framework. To our knowledge, the generalization error of the tilted empirical risk minimization has not been explored.
@@ 8 Conclusion and Future Work
Our current results are applicable to scenarios with bounded loss functions and in scenarios with negative tilting ($\gamma<0$) when the possibly unbounded loss function has a bounded second moment. However, our current results for the asymptotic regime $\gamma\rightarrow\infty$, where the tilted empirical risk is equal to maximum loss, are vacuous. Therefore, studying the generalization performance of tilted empirical risk minimization under unbounded loss functions for positive tilting, and obtaining informative bounds in the asymptotic regime $\gamma\rightarrow\infty$, are planned as future work. Moreover, we intend to apply other approaches from the literature, such as those discussed by Aminian et al. (2023), to derive further bounds on the tilted generalization error in the mean-field regime. As the tilted empirical risk has a better performance under the scenario of imbalanced samples for positive tilt ($\gamma>0$), it would be also interesting to study the tilted generalization error under imbalance scenarios.
@@ Acknowledgements
Appendix
[…]
Appendix A Other Related Works
This section details related works about the Rademacher complexity, stability and PAC-Bayesian bounds, as well as related work on unbounded loss functions.

### v2

@@ 4 Generalization Bounds for Unbounded Loss Functions
Under the same assumption in Theorem 4.3 and a finite hypothesis space, then for $n\geq\frac{(4\gamma^{2}\kappa_{u}^{2}+8/3\zeta)\log(2/\delta)}{\zeta^{2}\exp(2\gamma\kappa_{u})}$ with probability at least $(1-\delta)$ and $\gamma<0$, the excess risk of tilted empirical risk satisfies
$\mathfrak{E}_{\gamma}(\mu)$ $\leq\frac{4\kappa_{u}\exp(-\gamma\kappa_{u})}{(1-\zeta)}\sqrt{\frac{B(\delta)}{n}}-\gamma\kappa_{u}^{2}-\frac{8\exp(-\gamma\kappa_{u})B(\delta)}{3n\gamma(1-\zeta)},$ (22)
where $B(\delta)=\log(\mathrm{card}(\mathcal{H}))+\log(2/\delta)$.
@@ 5 Robustness of TERM
As shown in experiments by Li et al. (2021), the tilted empirical risk is robust to noise or outliers samples during training using negative tilt ($\gamma<0$). In this section, we study the robustness of the TER under distributional shift, $\tilde{\mu}$ and the following assumption is made.
@@
There is a constant $\kappa_{s}\in\mathbb{R}^{+}$ such that the loss function $(H,Z)\mapsto\ell(H,Z)$ satisfies $\mathbb{E}_{\tilde{\mu}}[\ell^{2}(h,Z)]\leq\kappa_{s}^{2}$ uniformly for all $h\in\mathcal{H}$.
@@
Given Assumption 5.1 and Assumption 4.1, then the difference of tilted population risk under, (9), between $\mu$ and $\tilde{\mu}$ is bounded as follows,
$\frac{1}{|\gamma|}\Big{|}\log(\mathbb{E}_{\tilde{Z}\sim\mu}[C(h,\tilde{Z})])-\log(\mathbb{E}_{\tilde{Z}\sim\tilde{\mu}}[C(h,\tilde{Z})])\Big{|}\leq\frac{\mathbb{TV}(\mu,\tilde{\mu})}{\gamma^{2}}\frac{\exp(|\gamma|\kappa_{u})-\exp(|\gamma|\kappa_{s})}{(\kappa_{u}-\kappa_{s})},$ (23)
where $C(h,\tilde{Z})=\exp(\gamma\ell(h,\tilde{Z}))$.
Note that, for positive $\gamma$, the result in Proposition 5.2 does not hold and can be unbounded. Using Proposition 5.2, we can provide an upper and lower bounds on the tilted generalization error under distributional shift.
Proposition 5.3.
Given Assumptions  4.1 and 5.1, for any fixed $h\in\mathcal{H}$ and with probability least $(1-\delta)$ for $\gamma<0$, then the following upper bound holds on the tilted generalization error
$\begin{split}&\mathrm{gen}_{\gamma}(h,\hat{S})\leq 2\kappa_{s}\exp(-\gamma\kappa_{s})\sqrt{\frac{\log(2/\delta)}{n}}-\frac{\gamma}{2}\kappa_{u}^{2}\\ &\quad-\frac{4\exp(-\gamma\kappa_{s})(\log(2/\delta))}{3n\gamma}+\frac{\mathbb{TV}(\mu,\tilde{\mu})}{\gamma^{2}}D(\kappa_{s},\kappa_{u}),\end{split}$
where $D(\kappa_{s},\kappa_{u})=\frac{\exp(|\gamma|\kappa_{u})-\exp(|\gamma|\kappa_{s})}{(\kappa_{u}-\kappa_{s})}$ and $\hat{S}$ is the training dataset under the distributional shift.
Proposition 5.4.
Given Assumptions  4.1 and 5.1, for any fixed $h\in\mathcal{H}$ and with probability least $(1-\delta)$, there exists a $\zeta\in(0,1)$ such that for $n\geq\frac{(4\gamma^{2}\kappa_{u}^{2}+8/3\zeta)\log(2/\delta)}{\zeta^{2}\exp(2\gamma\kappa_{u})}$ and $\gamma<0$, such that the following upper bound holds on the tilted generalization error
$\begin{split}&\mathrm{gen}_{\gamma}(h,\hat{S})\geq-\frac{2\kappa_{s}\exp(-\gamma\kappa_{s})}{(1-\zeta)}\sqrt{\frac{\log(2/\delta)}{n}}\\ &\qquad+\frac{4\exp(-\gamma\kappa_{s})(\log(2/\delta))}{3n\gamma(1-\zeta)}-\frac{\mathbb{TV}(\mu,\tilde{\mu})}{\gamma^{2}}D(\kappa_{s},\kappa_{u}),\end{split}$
where $D(\kappa_{s},\kappa_{u})=\frac{\exp(|\gamma|\kappa_{u})-\exp(|\gamma|\kappa_{s})}{(\kappa_{u}-\kappa_{s})}$ and $\hat{S}$ is the training dataset under the distributional shift.
Combining Proposition 5.3 and Proposition 5.4, we derive an upper bound on the absolute value of the tilted generalization error under distributional shift.
@@
Under the same assumptions in Proposition 5.4, then for $n\geq\frac{(4\gamma^{2}\kappa_{u}^{2}+8/3\zeta)\log(2/\delta)}{\zeta^{2}\exp(2\gamma\kappa_{u})}$ and $\gamma<0$, and with probability at least $(1-\delta)$, the absolute value of the tilted generalization error under distributional shift satisfies
$\begin{split}&\sup_{h\in\mathcal{H}}|\mathrm{gen}_{\gamma}(h,\hat{S})|\leq\frac{2\kappa_{s}\exp(-\gamma\kappa_{s})}{(1-\zeta)}\sqrt{\frac{B(\delta)}{n}}-\frac{4\exp(-\gamma\kappa_{s})B(\delta)}{3n\gamma(1-\zeta)}\\ &\qquad-\frac{\gamma\kappa_{u}^{2}}{2}+\frac{\mathbb{TV}(\mu,\tilde{\mu})}{\gamma^{2}}\frac{\big{(}\exp(|\gamma|\kappa_{u})-\exp(|\gamma|\kappa_{s})\big{)}}{(\kappa_{u}-\kappa_{s})},\end{split}$ (24)
where $B(\delta)=\log(\mathrm{card}(\mathcal{H}))+\log(2/\delta)$.
Similar results to Proposition 5.3, can be derived via an information-theoretic approach.
Robustness vs Generalization: The term $\frac{\mathbb{TV}(\mu,\tilde{\mu})}{\gamma^{2}}\frac{\big{(}\exp(|\gamma|\kappa_{u})-\exp(|\gamma|\kappa_{s})\big{)}}{(\kappa_{u}-\kappa_{s})}$ represents the distributional shift cost (or robustness) associated with the TER. This cost can be reduced by increasing $|\gamma|$. However, increasing $|\gamma|$ also amplifies other terms in the upper bound of the tilted generalization error. Therefore, there is a trade-off between robustness and generalization, particularly for $\gamma<0$ in the TER. Interestingly, Li et al. (2021) also observed this trade-off for negative tilt.
@@ 6 The KL-Regularized TERM Problem
where $I(H;S)\leq\mathrm{KL}(P_{H|S}\|\pi_{H}|P_{S})$ and $\pi_{H}$ is a prior distribution over hypothesis space $\mathcal{H}$. All proof details are deferred to Appendix G.
@@
The solution to the expected TERM regularized via KL divergence, (26), is the tilted Gibbs Posterior (a.k.a. Gibbs Algorithm),
@@
is the solution to the KL-regularized ERM minimization problem, where $\tilde{F}_{\alpha}(S)$ is the normalization factor. Therefore, the tilted Gibbs posterior is different from the Gibbs posterior, (28). It can be shown that for $\gamma\rightarrow 0$, the tilted Gibbs posterior converges to the Gibbs posterior. Therefore, it is interesting to study the expected generalization error of the tilted Gibbs posterior. For this purpose, we give an exact characterization of the difference between the expected TER under the joint and the product of marginal distributions of $H$ and $S$. More results regarding the Gibbs posterior are provided in Appendix G.
@@
$\begin{split}&\overline{\mathrm{gen}}_{\gamma}(H,S)\leq\frac{\alpha(\exp(|\gamma|M)-1)^{2}}{2\gamma^{2}n}\\ &-\frac{\gamma\exp(-\gamma M)}{2}\big{(}1-\frac{1}{n}\big{)}\mathbb{E}_{P_{H}}\big{[}\mathrm{Var}_{\tilde{Z}\sim\mu}(\ell(H,\tilde{Z}))\big{]}.\end{split}$ (30)
@@
$\begin{split}&|\overline{\mathrm{gen}}_{\gamma}(H,S)|\leq\frac{\alpha(\exp(|\gamma|M)-1)^{2}}{2\gamma^{2}n}\\ &+\frac{|\gamma|M^{2}\exp(|\gamma|M)}{8}\Big{(}1-\frac{1}{n}\Big{)}.\end{split}$ (31)
@@ 7 Related Works
Generalization Error Analysis: Different approaches have been applied to study the generalization error of general learning problems under empirical risk minimization, including VC dimension-based, Rademacher complexity, PAC-Bayesian, stability and information-theoretic bounds. In this section, we discuss the related works about Uniform and information-theoretic bounds. More related works for generalization error analysis are discussed in Appendix B.
@@
The aforementioned approaches are applied to study the generalization error in the linear empirical risk framework. To our knowledge, the generalization error of the tilted empirical risk minimization from information theoretical and uniform approach perspectives has not been explored.
@@ 8 Conclusion and Future Work
Our current results are applicable to scenarios with bounded loss functions and in scenarios with negative tilting ($\gamma<0$) when the possibly unbounded loss function has a bounded second moment. However, our current results for the asymptotic regime $\gamma\rightarrow\infty$, where the tilted empirical risk is equal to maximum loss, are vacuous. Therefore, studying the generalization performance of tilted empirical risk minimization under unbounded loss functions for positive tilting, and obtaining informative bounds in the asymptotic regime $\gamma\rightarrow\infty$, are planned as future work. Moreover, we intend to apply other approaches from the literature, such as those discussed by Aminian et al. (2023), to derive further bounds on the tilted generalization error in the mean-field regime.
@@ Acknowledgements
Appendix A Overview of Main Results
An overview of our main results is provided in Fig. 1. All notations are summarized in Table 1.
Table 1: Summary of notations in the paper
Notation Definition Notation Definition
$S$ Training set $Z_{i}$ $i$-th sample
$\mathcal{X}$ Input (feature) space $\mathcal{Y}$ Output (label) space
$\mu$ Data-generating distribution $\tilde{\mu}$ Data-generating distribution under distributional shift
$\gamma$ Tilt parameter $\mathrm{KL}(P\|Q)$ KL-divergence between $P$ and $Q$
$I(X;Y)$ Mutual information $D_{\mathrm{SKL}}(P\|Q)$ Symmetrized KL divergence
$\mathbb{TV}(P,Q)$ Total variation distance $\kappa_{u}$ Bound on second moment
$h$ Hypothesis $\mathcal{H}$ Hypothesis space
$\ell(h,z)$ Loss function $R(h,\mu)$ Population (true) risk
$\widehat{\mathrm{R}}_{\gamma}(h,S)$ Tilted empirical risk $\mathrm{gen}_{\gamma}(h,S)$ Tilted generalization error
$\overline{\mathrm{R}}_{\gamma}(h,P_{S})$ Tilted population risk $\mathfrak{E}_{\gamma}(\mu)$ Excess risk
Tilted Generalization Error Bounded Loss Function ($\gamma\in\mathbb{R}$) Section 3 Information-theoretic Bound Upper bound (Theorem 3.13) & Lower bound (Theorem 3.14) Uniform Bound Upper bound (Theorem 3.3) & Lower bound (Theorem 3.4) Unbounded Loss Function ($\gamma\in\mathbb{R}_{<0}$) Section 4 Information-theoretic Bound Upper bound (Theorem E.4) & Lower bound (Theorem E.6) Uniform Bound Upper bound (Theorem 4.2) & Lower bound (Theorem 4.3) Robustness of TERM ($\gamma\in\mathbb{R}_{<0}$) Section 5 Uniform Bound Upper bound (Proposition 5.3) & Lower bound (Theorem 5.4) KL-Regularized TERM ($\gamma\in\mathbb{R}$) Section 6 Information-theoretic Bound Upper bound (Theorem 6.3)
Figure 1: Overview of the main results
Appendix B Other Related Works
This section details related works about Tilted empirical risk minimization, the Rademacher complexity, stability and PAC-Bayesian bounds, as well as related work on unbounded loss functions.